# Generating Temporal Networks with the Ascona Model

### v1

@@ 3.1.3 Conflicts
A conflict arises in the sampling mechanism whenever we want to resample an already present link. A possible cause for such type of conflicts is that the queue is congested and close to saturation. We can estimate if a model is prone to this problem by verifying its likelihood of getting close to the saturation value $\binom{N}{2}$. If a queue has a stationary value much lower than that value, and the variance of its trajectories is limited, this likelihood will be small, and conflicts due to congestion will occur rarely.
@@
Another cause for conflicts is connectivity constraints. For instance, consider the situation of a node $v_{i}$ already connected to a node $v_{j}$, and a new link originates from $v_{i}$. If, according to the probabilistic connectivity matrix, $v_{i}$ connects exclusively with a node $v_{j}$, the new link will generate a conflict. The number of issues hence increases if the connectivity constraints increase the probability of resampling the same pairs. We can gauge how problematic a certain connectivity probability matrix is by computing its Shannon entropy. Similarly, to gauge a connectivity conditional probability matrix we can compute its conditional entropy conditioned on the sampling vector; analogously to [koovely_evolution_2025], where the probabilistic matrix is a transition operator between nodes instead of a connectivity one.
If a conflict arises, it means that after a link $(v_{1},v_{2},s_{1},e_{1})$, we sampled a link $(v_{1},v_{2},s_{2},e_{2})$ with $s_{2}<e_{2}$. Three possible solutions are:
  • merge the two links in one $(v_{1},v_{2},s_{1},\max(e_{1},e_{2}))$,
  • discard all elements of the second link,
  • resample the pair of nodes but keep the interval $(s_{2},e_{2})$ as basis for the link.
All suggested solutions lead to small discrepancies from the desired temporal or connectivity structure. One might choose a solution to mitigate discrepancies from one or the other aspect, or based on practical considerations: for instance, if a queue is saturated, resampling nodes won’t work.
Notice that the likelihood of errors decreases as the number of nodes increases. The presence of an error does not increase the probability of another error occurring, so they do not propagate, and actually, there is a self-adjustment mechanism: if the appearance of a link causes a conflict, then the discrepancy caused by it ends before the problematic link’s disappearance.
@@
Figure 2: Example of queue blocks $L^{\prime}$ and $L^{\prime\prime}$ sharing the same links’ repartition in time (sampled from a $M/M/\infty$ process), but different connectivity structures. For each of the two sampled link streams, we display the adjacency matrix of the instantaneous graphs $G_{t}^{\prime}$ and $G_{t}^{\prime\prime}$ at three different times $(t\in\{20,50,80\})$. On top, we show the ones of $L^{\prime}$, that has a constant two-block connectivity structure. On the bottom side, we show the ones of $L^{\prime\prime}$, that does not have any particular structure: all pairs of nodes have the same probability to be connected, hence this could be considered a link stream version of an ER graph.
3.2 EDEDE Structures
@@
An EDEDE structure is a link stream generated based on the following inputs: a starting time $t_{s}$, an ending time $t_{e}$, an activity rate $\lambda$, and an interaction rate $\mu$. First, we impose that the starting times of each event are distanced by i.i.d. exponential random variables with parameter $\lambda$, or equivalently that the starting times are Poisson distributed with average $\lambda(t_{e}-t_{s})$. Additionally, we require that all events have i.i.d. exponential duration with parameter $\mu$.
We therefore have Exponential-Duration Events Distanced Exponentially, hence the acronym EDEDE.
@@
We can conveniently sample such structures with the Ascona method. The underlying queue model is $M/M/\binom{N}{2}/\binom{N}{2}/\infty/\text{FCFS}$ [gross_fundamentals_2008, Ch. 2.8], a truncated version of the $M/M/\infty$ queue [robert_stochastic_2013, Ch. 6.7]. However, we can assume to be working in the regime $\binom{N}{2}\gg\lambda\mu$ so that we can consider the distortions due the finite-source constraint as boundary effects and the model is approximately a $M/M/\infty$ model.
@@
Figure 3: Representation of the generation of an EDEDE structure with the Ascona model. (A) The first step consists of drawing i.i.d. exponentially distributed time intervals with rate $\lambda$, or conceptually equivalently, time points based on a Poisson process with rate $\lambda$, over a desired time interval. Each sampled time point corresponds to a starting time $s_{i}$ of a link of the dynamic network; in this example, there are five of them: $s_{1},\dots,s_{5}$. (B) The second step consists of sampling a time interval for each drawn starting point based on i.i.d. exponentially distributed time intervals with rate $\mu$. Each drawn interval describes the duration of the corresponding link. In this example, we end up with five link intervals: $(s_{1},e_{1}),\cdots,(s_{5},e_{5})$. (C) Finally, each link is associated with a pair of vertices. For each link, the first element of the pairs is drawn uniformly at random from the node set. Its interacting partner is drawn randomly based on a distribution encoded in the connectivity conditional probability matrix displayed on the left. In this case, the connectivity matrix encodes a two-block structure with high probability of generating links within each block, and low probability of link generation between blocks. We display node names on a vertical axis on the left of the figure. Each link is represented by a coloured rectangle showing the involved nodes and the time interval when it happens.
@@
From an information-theoretical or Bayesian perspective, an EDEDE structure can be seen as a reference null model that does not assume any knowledge of any correlation between arrival and service times: the exponential distributions are uninformative priors because they have maximum differential entropy among all continuous distributions supported in $0,\infty]$ with the same mean [dowson_maximum-entropy_1973].
@@
However, one might be interested in sampling links based on different distributions, other than exponential ones [barucca_maximum_2025]. In that case, we obtain queues alternative to the $M/M/\infty$ one, whose implementation leads to small modifications of the Ascona sampling method, but lead to structures different from the EDEDE ones. Theoretical properties of these derived models depend on the theoretical properties of these alternative queueing processes; however, some properties like the stationary value are often independent from the Markovian assumptions [bunday_gmr_1980, gross_fundamentals_2008]. See [gross_fundamentals_2008] for a discussion of the consequences of dropping the purely Markovian assumption for the arrival times, in particular allowing restrictions on the support, inhomogeneity of the rate parameter, and batch arrivals.
@@ 3.2.1 Degenerate Cases
We can generate link streams with 0-duration links with the Ascona model by setting the service times to be Dirac $\delta$ distributed. In the EDEDE structure case, this reduces to a system with links spawning as a Poisson process.
@@
3.3 Smooth Stochastic Blocks
In this part, we describe how to use the Ascona model to sample link streams with evolving community structures. The idea is to design a single queue block where the connectivity follows the desired clusters’ evolution. In this article, we show EDEDE samples exclusively, but the desired effect is achievable almost independently of the underlying queue process by imposing an appropriate connectivity (conditional) probability matrix that evolves as a function of time.
In the example displayed in Figure˜2, we used constant two-block connectivity conditional probability matrices, leading to a constant community structure. In the example displayed in Figure˜5, we first have a four-block structure followed by a two-block structure. The change represents a discontinuity in the progression of the underlying connectivity conditional probability matrix, however preexisting links take a while to disappear after the change, while links associated with the new structure appear gradually, leading to smooth transitions, hence the name “smooth stochastic blocks".
In Section˜3.5, we explain how to aggregate smooth stochastic blocks in order to get snapshots, and after that, we compare them to Dynamic SBMs, the main generalization of SBMs to temporal networks. Now, we showcase the expressivity of the newly introduced smooth SBMs.
3.4 Generating Archetpyes
A fundamental aspect of temporal networks is the understanding of its clusters and how they evolve in time. Similarly to the methods used in [asgari_mosaic_2023] and [cazabet_evaluating_2021], the procedure to generate link stream scenarios follows two steps: first, the experimenter needs to describe the scenario in terms of a queue blocks’ footprint, and secondly, generate the structure with the Ascona model to obtain a temporal network that qualitatively reflects the desired scenario.
In this part, we display the expressivity of the Ascona model by describing how to generate the typical “events" characterizing temporal networks’ evolution listed in [cazabet_challenges_2023]. For elementary versions of such scenarios, a single EDEDE queue block is sufficient. In order to design some of these events, one can focus exclusively on the temporal aspect of the model.
@@ 3.4 Generating Archetpyes
In a birth event, a new community appears. Conversely, in a death event, an existing community vanishes. One can model the birth of a community by either taking the head or the time-reversed tail of an EDEDE queue block. In the first case, the growth is not necessarily monotonic and has roughly a concave shape. In the second case, the growth is monotonic and has approximately a convex shape. Conversely, one can model the death of a community by taking the time-reverse of a birth event.
@@
One event, not listed among the ones in [cazabet_challenges_2023], is the one where a stable community experiences an increase/decrease in the number of interactions. We can model this situation with an EDEDE queue block by imposing that the parameters $\lambda$ and $\mu$ of the underlying $M/M/\infty$ process change at a switching time point $t_{sw}$. The underlying process remains Poissonian, and its average is
@@
More complex temporal scenarios require a chain of queue blocks.
@@
Resurgency can be thought of as a combined event where a queue block vanishes for a period, then reappears without particular perturbations.
@@
The growth and contraction events are similar to the previous case. In a growth event, we observe an existing community acquire new nodes; in a contraction, a community loses some nodes, but does not disappear completely. In these cases, instead of two interacting communities, one community interacts with a set of unstructured, lowly connected nodes.
One can notice that this procedure leads to controllably smooth transitions, balancing stability in time of the links with consistency with the connectivity structures, a known problematic aspect of sampling temporal networks [cazabet_evaluating_2021, Section 3.2.1].
We display in Figure˜5 the generation of many of these events with a single EDEDE block, encoding the evolution of a link stream from a four-block to a two-block structure. We show as well that by slicing the time domain and aggregating the information over the slices, we get a snapshot equivalent of this link stream. We discuss this procedure in more detail in the next part.
Figure 5: Ascona model sample displaying various events. (a) Birth event of four communities. (b) In the stationary regime, there is a stable configuration with four fully developed communities. (c) At time 100, there is a merge event leading to the formation of two bigger communities. (d) After time 200, the queue decays, and the two communities vanish. (A) Weighted footprint of the network of the interval $[40s,60s]$ displaying clearly the four-block structure before the merge. (B) Weighted footprint of the interval $[90s,110s]$ displaying the transition between the four and the two-block structure happening at time $100s$. (C) Weighted footprint of the interval $[140s,160s]$ displaying the two-block structure after the merge. The colour of each footprint entry depends on the weight from Eq. (2.4). Intuitively, warmer colours indicate higher weights.
@@ 3.5 Generating Snapshot Networks
As mentioned previously, the majority of datasets and methods in temporal networks theory has been designed primarily (and often exclusively) with snapshot networks in mind. We therefore describe a procedure for generating snapshot networks from an EDEDE sample in a controlled way. The strategy is to discard the head of the queue based on the estimated crossover times (Eq. (2.28) or its approximation Eq. (2.29)) presented in Section˜2.2.3, and then slice the time domain and aggregate the links in unweighted footprints (see Eq. (2.3)) which leads to “uniform" snapshots.
@@
In both cases, one could increase the overlap between subsequent slices to achieve a higher level of smoothness. Notice that if one wants to exert control on the number of distinct links in every slice, then a correction factor for Eq. (3.3) and Eq. (3.5) is needed to take into account the probability of links appearing multiple times in a certain $\Delta t$.
@@ 3.5.1 Comparison with Dynamic SBM
If we apply the above-presented procedure to smooth SBM link streams, we obtain snapshots with block structures. This is reminiscent of the Dynamic SBM model, which has emerged as a popular extension of its static network counterpart and is used, for instance, for testing change-point detection methods (e.g., in [bhattacharjee_change_2020, sulem_graph_2024]). In dynamic SBMs, each snapshot is sampled according to a static SBM distribution whose parametrization can change in time. In terms of expected number of edges and its variance, smooth SBMs and dynamic SBMs are similar because of the similarity between Poisson and Binomial distributions. The two methods differ in terms of smoothness: dynamic SBM snapshots have no mechanism to enforce smoothness, while the smoothness of the underlying smooth SBM link stream is inherited by the derived SBM snapshots.

### v2

@@ 3.1.4 Conflicts
A conflict arises whenever we attempt to add a link that overlaps in time with an already present link on the same unordered pair. Concretely, a proposed link $\ell_{2}=([s_{2},e_{2}),\{u,v\})$ conflicts with an existing link $\ell_{1}=([s_{1},e_{1}),\{u,v\})$ if $s_{2}<e_{1}$ and $s_{1}<e_{2}$.
If multi-links are allowed (i.e. $E$ is treated as a multiset), the sampler draws exactly from the intended construction. If we enforce simplicity (at most one active link per pair at any time), conflict handling can be interpreted as a projection from an ideal multiset-valued sample to the space of simple link streams, and the sampler becomes approximate.
When looking at a single queue block, there are two main causes for conflicts:
Congestion / saturation
A simple link stream on $N$ nodes admits at most $\binom{N}{2}$ simultaneous active links. The temporal mechanism may still be specified by a queue with an abstract capacity $K>\binom{N}{2}$, but then enforcing simplicity requires a projection step. We can estimate if a block is prone to this problem by looking at the typical occupancy level of the underlying queue: if it is well below $\binom{N}{2}$, such projections are rarely activated.
Endpoint concentration
@@
Another cause for conflicts is due to connectivity constraints. For instance, if a node $v_{i}$ is already connected to $v_{j}$ and the endpoint law forces new links out of $v_{i}$ to connect almost exclusively to $v_{j}$, then a new proposed interval is likely to target $\{v_{i},v_{j}\}$ again. More generally, if the endpoint law heavily favors a small subset of pairs, resampling an already active pair becomes likely even far from saturation. We can gauge how problematic a certain connectivity probability matrix is by computing its Shannon entropy. Similarly, to gauge a connectivity conditional probability matrix, we can compute its conditional entropy conditioned on the sampling vector; analogously to [19], where the probabilistic matrix is a transition operator between nodes rather than a connectivity one.
Possible projection rules include:
  • Merge: replace $\ell_{1}$ and $\ell_{2}$ by $([\min(s_{1},s_{2}),\max(e_{1},e_{2})),\{u,v\})$,
  • Discard: ignore $\ell_{2}$ and keep $\ell_{1}$,
  • Resample endpoints: keep the interval $s_{2},e_{2})$ but resample the unordered pair $\{u,v\}$.
Each rule introduces small, local deviations from the target temporal/connectivity laws. One might choose a rule to mitigate discrepancies from one or the other aspect, or based on practical considerations: for instance, if a queue is saturated, resampling nodes won’t help. Importantly, the presence of an error does not increase the probability of another error occurring, so deviations do not propagate, and actually, there is a self-adjustment mechanism: once the overlapping interval between conflicting links ends, the discrepancy disappears.
Superposition of blocks can cause conflicts as well and can be resolved with the same projection rules. We summarize the sampling procedure in Algorithm˜1.
Algorithm 1 Ascona sampler (block-based)
0: Vertex set $V$; blocks $(\mathsf{Queue}_{r},\ t_{e}^{(r)},\ a_{r},\ \text{trim}_{r},\ \text{endpoint-law}_{r})_{r=1}^{R}$; conflict rule $\mathsf{Resolve}$
0: Link stream $L=(T,V,E)$
[…]
2:for $r=1$ to $R$ do
3:  Sample intervals $\mathcal{I}_{r}=\{[T_{k},E_{k})\}_{k=1}^{n_{r}}$ from $\mathsf{Queue}_{r}$ on $0,t_{e}^{(r)}]$, apply $\text{trim}_{r}$
4:  for each $T_{k},E_{k})\in\mathcal{I}_{r}$ do
5:   Sample unordered pair $\{U_{k},V_{k}\}$ from $\text{endpoint-law}_{r}$ {independent of $\mathcal{I}_{r}$}
6:   Propose link $\ell=([T_{k}+a_{r},E_{k}+a_{r}),\{U_{k},V_{k}\})$
7:   $E\leftarrow\mathsf{Resolve}(E,\ell)$ {projection step}
8:  end for
9:end for
[…]
11:return $L=(T,V,E)$
We show a concrete example of two dynamic networks sampled with the Ascona sampler in Figure˜2. The two samples share the same temporal structure but differ in their endpoint laws. The shared temporal structure is sampled based on an $M/M/\infty$ queue. We now discuss in more detail this fundamental parametrization, whose construction is illustrated in Figure˜3.
@@
Figure 2: Example of queue blocks $L^{\prime}$ and $L^{\prime\prime}$ sharing the same links’ repartition in time (sampled from a $M/M/\infty$ process), but different connectivity structures. For each of the two sampled link streams, we display the adjacency matrix of the instantaneous graphs $G_{t}^{\prime}$ and $G_{t}^{\prime\prime}$ at three different times $(t\in\{20,50,80\})$. On top, we show the ones of $L^{\prime}$, which has a constant two-block connectivity structure. On the bottom side, we show the ones of $L^{\prime\prime}$, which does not have any particular structure: all pairs of nodes have the same probability to be connected, hence this could be considered a link stream version of an ER graph.
3.2 EDLDE Structures
@@
An EDLDE structure is a link stream generated based on the following inputs: a starting time $t_{s}$, an ending time $t_{e}$, an activity rate $\lambda$, and an interaction rate $\mu$. We assume that link start times form a homogeneous Poisson process on $t_{s},t_{e}]$ with rate $\lambda$. Equivalently, the number of starts in $t_{s},t_{e}]$ is $\mathrm{Poisson}(\lambda(t_{e}-t_{s}))$. Additionally, link durations are i.i.d. exponential with rate $\mu$. We therefore have Exponential-Duration Links Distanced Exponentially, hence the acronym EDLDE.
@@
We can conveniently sample such structures with the Ascona method. The underlying temporal mechanism corresponds to an $M/M/\infty$ queue, optionally capped, as in $M/M/\binom{N}{2}/\binom{N}{2}/\infty/\mathrm{FCFS}$ [16, Ch. 2.8], a truncated counterpart of $M/M/\infty$ [27, Ch. 6.7]. In the regime $\binom{N}{2}\gg\frac{\lambda}{\mu},$ the occupancy is far from saturation, so distortions due to the finite-pair constraint are boundary effects and the temporal mechanism is well approximated by the $M/M/\infty$ queue.
@@
Figure 3: Representation of the generation of an EDLDE structure with the Ascona model. (A) The first step consists of drawing i.i.d. exponentially distributed inter-arrival times with rate $\lambda$, or equivalently sampling start times from a Poisson process with rate $\lambda$, on a desired time interval. Each sampled time corresponds to a starting time $T_{i}$ of a link; in this example, there are five of them: $T_{1},\dots,T_{5}$. (B) The second step consists of sampling an i.i.d. exponential duration with rate $\mu$ for each start time. In this example, we obtain five link intervals $T_{1},E_{1}),\dots,[T_{5},E_{5})$. (C) Finally, each interval is assigned an unordered pair of vertices. For each link, the first endpoint is drawn uniformly at random from the node set and the partner is drawn from the connectivity conditional probability matrix displayed on the left. In this case, the matrix encodes a two-block structure with high within-block probability and low between-block probability. We display node names on a vertical axis on the left of the figure. Each link is represented by a colored rectangle showing the involved nodes and the time interval when it occurs.
@@
From an information-theoretical or Bayesian perspective, an EDLDE structure can be seen as a reference null model that does not assume any knowledge of any correlation between arrival and service times: exponential distributions are uninformative priors because they have maximum differential entropy among all continuous distributions supported on $0,\infty)$ with the same mean [12].
@@
However, one might be interested in sampling links based on other distributions than exponential ones. In that case, we obtain queues alternative to the $M/M/\infty$ one. Their implementation leads to small modifications of the Ascona sampling method and produces structures different from the EDLDE ones. Theoretical properties of these derived models depend on the corresponding properties of the alternative queueing processes; however, some properties, like the stationary mean occupancy, are often independent of the purely Markovian assumptions [7, 16]. See [16] for a discussion of the consequences of dropping the Markovian assumption for the arrival process, in particular allowing restrictions on the support, inhomogeneity of the rate parameter, and batch arrivals.
@@ 3.2.1 Degenerate Cases
We can generate link streams with 0-duration links with the Ascona model by setting the service times to be Dirac $\delta$ distributed. In the EDLDE structure case, this reduces to a system with links spawning as a Poisson process.
@@
3.2.2 Smooth Stochastic Blocks
In this part, we describe how to use the Ascona model to sample link streams with evolving community structures. The idea is to concatenate blocks, or equivalently, to design a single queue block whose connectivity (conditional) probability matrix evolves as a function of time and follows the desired clusters’ evolution. In this article, we show EDLDE samples exclusively, but the desired effect is achievable almost independently of the underlying queue.
In the example displayed in Figure˜2, we used constant two-block connectivity conditional probability matrices, leading to a constant community structure. In the example displayed in Figure˜5, we first have a four-block structure followed by a two-block structure. The change represents a discontinuity in the progression of the underlying connectivity conditional probability matrix; however, preexisting links take a while to disappear after the change, while links associated with the new structure appear gradually, leading to smooth transitions, hence the name “smooth stochastic blocks".
In Section˜3.3, we explain how to aggregate smooth stochastic blocks in order to get snapshots, and after that, we compare them to Dynamic SBMs, the main generalization of SBMs to temporal networks. Now, we showcase the expressivity of the newly introduced smooth SBMs.
3.2.3 Generating Archetypes
A fundamental aspect of temporal networks is the understanding of their clusters and how they evolve in time. Similarly to the methods used in [2] and [9], the procedure to generate link stream scenarios follows two steps: first, the experimenter needs to describe the scenario in terms of a queue block’s footprint, and secondly, generate the structure with the Ascona model to obtain a temporal network that qualitatively reflects the desired scenario.
In this part, we display the expressivity of the Ascona model by describing how to generate the typical “events" characterizing temporal networks’ evolution listed in [10]. For elementary versions of such scenarios, a single EDLDE queue block is sufficient. In order to design some of these events, one can focus exclusively on the temporal aspect of the model.
@@ 3.2.3 Generating Archetypes
In a birth event, a new community appears. Conversely, in a death event, an existing community vanishes. One can model the birth of a community by either taking the head or the time-reversed tail of an EDLDE queue block. In the first case, the growth is not necessarily monotonic and has roughly a concave shape. In the second case, the growth is monotonic and has approximately a convex shape. Conversely, one can model the death of a community by taking the time-reverse of a birth event.
@@
One event, not listed among the ones in [10], is the one where a stable community experiences an increase/decrease in the number of interactions. We can model this situation with an EDLDE queue block by imposing that the parameters $\lambda$ and $\mu$ of the underlying $M/M/\infty$ process change at a switching time point $t_{sw}$. The underlying process remains Poissonian, and its average is
@@
More complex temporal scenarios require more complex concatenations.
@@
Resurgence can be thought of as a combined event where a queue block vanishes for a period, then reappears without particular perturbations.
@@
The growth and contraction events are similar to the previous case. In a growth event, we observe an existing community acquires new nodes; in a contraction, a community loses some nodes, but does not disappear completely. In these cases, instead of two interacting communities, one community interacts with a set of unstructured, lowly connected nodes.
One can notice that this procedure leads to controllably smooth transitions, balancing stability in time of the links with consistency with the connectivity structures, a known problematic aspect of sampling temporal networks [9, Section 3.2.1].
We display in Figure˜5 the generation of many of these events with a single EDLDE block, encoding the evolution of a link stream from a four-block to a two-block structure. We show as well that by slicing the time domain and aggregating the information over the slices, we get a snapshot equivalent of this link stream. We discuss this procedure in more detail in the next part.
Figure 5: Ascona model sample displaying various events. (a) Birth event of four communities. (b) In the stationary regime, there is a stable configuration with four fully developed communities. (c) At time 100, there is a merge event leading to the formation of two bigger communities. (d) After time 200, the queue decays, and the two communities vanish. (A) Weighted footprint of the network of the interval [40s,60s][40\,s,60\,s] displaying clearly the four-block structure before the merge. (B) Weighted footprint of the interval $[90s,110s]$ displaying the transition between the four and the two-block structure happening at time $100s$. (C) Weighted footprint of the interval $[140s,160s]$ displaying the two-block structure after the merge. The colour of each footprint entry depends on the weight from eq. (2.4). Intuitively, warmer colours indicate higher weights.
@@ 3.3 Generating Snapshot Networks
As mentioned previously, the majority of datasets and methods in temporal networks theory have been designed primarily (and often exclusively) with snapshot networks in mind. We therefore describe a procedure for generating snapshot networks from an EDLDE sample in a controlled way. The strategy is to discard the head of the queue based on the estimated crossover times (eq. (2.28) or its approximation eq. (2.29)) presented in Section˜2.2.3, and then slice the time domain and aggregate the links in unweighted footprints (see eq. (2.3)), which leads to “uniform" snapshots.
@@
In both cases, one could increase the overlap between subsequent slices to achieve a higher level of smoothness. Notice that if one wants to exert control on the number of distinct links in every slice, then a correction factor for eq. (3.10) and eq. (3.12) is required to take into account the probability of links appearing multiple times in a certain $\Delta t$.
@@ 3.3.1 Comparison with Dynamic SBM
If we apply the above-presented procedure to smooth SBM link streams, we obtain snapshots with block structures. This is reminiscent of the Dynamic SBM model, which has emerged as a popular extension of its static network counterpart and is used, for instance, for testing change-point detection methods (e.g., in [5, 33]). In dynamic SBMs, each snapshot is sampled according to a static SBM distribution whose parametrization can change in time. In terms of expected number of edges and its variance, smooth SBMs and dynamic SBMs are similar because of the similarity between Poisson and Binomial distributions. The two methods differ in terms of smoothness: dynamic SBM snapshots have no mechanism to enforce smoothness, while the smoothness of the underlying smooth SBM link stream is inherited by the derived SBM snapshots.